t varies directly as the square root of l. If l=81, then t=10 what is the constant proportion?
1. a is directly proportional to b. If a = 15, then b = 9
2. M varies directly as n. If n = 2/3, then m = 1/4
3. T varies as the square root of L. If L = 81, then T = 10
4. Z is proportional to the cube of d. If d = 2 then z = 5
5. f varies directly as the square of t. If t = 16, then r = 6
6. p varies directly as the sum of L and W. If L = 4 and W = 3, then P = 14
7. C is directly proportional to r. If r = 5, then c = 10(pi)
8. Y varies directly as the cube root of x. If x = 27, then y = 2z
9. J varies directly as E. If J = 594, then E = 225.
10. X varies as 1/y. If X = 25, y = 10
Answer #1
Wow , looks like you want us to do your assignment for you.

I will do one of them, say #8, then you do the rest

#8
Y varies directly as the cube root of x
y = k (x)^(1/3)
given : when x=27, y = 2z
2z = k(27)^(1/3)
2z = k(3)
k = 2z/3

(was this a typo? all the other cases had numbers)
answered by: Veil